This CD contains the freely distributable version of the OpenSSH client for various Unixes and its associated file transfer program, scp and sftp. It also contains the corresponding programs for M$ Windows and Mac OS.

On the CD there are files README (for Unix users), README.TXT (For Windows users) and README.MAC (for MacOS users). These files have more details of the CD content and how to use it. On-line documentation may

be found at http://www-uxsup.csx.cam.ac.uk/CD/ .

On Solaris, Irix, Windows and MacOS a window showing the CD’s content should appear automatically on insertion of the CD.

On a Debian system try “cd /var/autofs/misc/” and looking. If the autofs package is installed then the OS will automount the CD there. On Red Hat Linux try running the command “mount /mnt/cdrom” when logged on at the console. If that works then you will need to do “umount /mnt/cdrom” when you wish to eject the CD.

If none of the above works you will have to contact the system’s administrator to mount the CD.

Queries by e-mail only, please, to: unix-support@ucs.cam.ac.uk

CD id: UCAM SSH CD 05a